In a thunderstorm, lightning can transfer more tha
Answer #1

Lets calculate the capacitance of cloud,

C = eo A/d =8.85*10-12*pi*1100*1100/1600 = 2.1*10-8 F

now charge, Q=CV

Q=1020*1.6*10-19 Because charge of 1 electron is1.6*10-19

so potential difference, V=Q/C = 1020*1.6*10-19/[2.1*10-8 ] = 7.77* 10-9 V